A HISTORY OF THE BANACH SPACE ARCHIVE

AND

IMPLICATIONS FOR ELECTRONIC ARCHIVES OF

PUBLICATIONS (PRELIMINARY VERSION)

DALE ALSPACH

The Banach space list and archive was begun in August 1989 and was founded using a list serving package on a DEC VMS VAX. I became aware of this type of software for the first time and realized that it had the potential to implement an idea for a Banach space newsletter that Pete Casazza had proposed. Pete had been discussing the idea of having a newsletter that appeared several times a year and would announce new results, advertise conferences, and provide a forum for queries. Because of the amount of work, startup costs, and other considerations Pete’s idea was never implemented in paper form. The advent of email and list server software made it possible to implement some of Pete’s ideas without the costs of production and of mailing a printed newsletter.

The initial list originated from electronic address lists that had been gathered from individual researchers in Banach space theory. At that time there were between 30 and 40 addresses that were used regularly and it was common to receive a message with all of the addresses listed. This list became the subscriber list at start up of the Banach archive. The list functioned by email only. I had no real experience with running such a list and no experience with being a subscriber to one. Thus the way the list functioned evolved slowly as I gained experience.

At first the list was unmoderated. Anyone who knew the address could send something to everyone on the list. Thus often papers that were in TeX or LaTeX were sent to all subscribers. For a time this was satisfactory, however several circumstances led to a change. One problem was the repeated accidental distribution of messages. In 1989 email was a new thing for most people and many of the people using it were not very familiar with computers. It seems to be an initiation rite for almost everyone who uses a list to confuse the list command address with the distribution address and unintentionally send a subscription

Date: January 8, 1999.

command to all of the list subscribers. This happened frequently in the early days of the Banach archive and I have noted that it still happens on many unmoderated lists today. This is an annoyance but there were also some accidental distributions of sensitive data. Finally some users complained that because of cost and/or limitations on mail inbox size that large mail messages were unwelcome. (The practice of charging for storage space or bytes transmitted no longer affects most of us but may return someday. [Metc, MM]) The list thus became moderated and whenever a paper was added to the archive, only the abstract was emailed to all subscribers. As the number of papers added to the archive increased so did my workload and I began to look for ways to enhance the functionality of the archive and decrease the demands on me. In order to provide uniform information about each paper and to automate some of the pro- cedures involved with posting papers, a required header for each paper was designed and instituted. This included obvious items such as the authors’ names, the title, an abstract, a mathematics subject classifi- cation, TeX format, etc. One item that presently would seem unusual that was included on the header was a list of the standard printable characters and their decimal ASCII codes (32–126). At the time the Banach archive began there were some standardization problems with the translation of certain characters between machines. The principal problem was between EBCDIC machines such as many IBM machines and those following the ASCII standard. The problem was particularly disastrous for TeX files because braces were often incorrectly translated. By examining this part of the header one could quickly determine if the email had passed through an incorrect translation and thus was damaged. %Special character check block %32 space 33 ! exclam. pt. 34 " double quote 35 # sharp

%36 $ dollar       37 % percent       38 & ampersand     39 ' prime

%40 ( left paren.  41 ) rt. paren.    42 * asterisk      43 + plus

%44 , comma        45 - minus         46 . period        47 / divide

%58 : colon        59 ; semi-colon    60 < less than     61 = equal

%62 > greater than 63 ? question mark 64 @ at

%91 [ left bracket 92 \ backslash     93 ] right bracket 94 ^caret

%95 _underline    96 ` left single quote

%123 { left brace  124 | vertical bar 125 } right brace  126 ~ tilda

The moderation of the bulletin board prevented accidental mailings, but it had its own problems. One question that arose was what messages were appropriate to distribute to the subscribers to a Banach space list. At first I screened only for offensive material (of which there